A payphone service charges php 5 for the first three minutes and php 1 for every additional or a fraction of it. How much will the caller have to pay if his call last 8 minutes. Write a rule the best describes the problem and draw its graph using an method
Answer by ikleyn(52931)

.

    How much = 5 + 1*(8-3) = 5 + 5 = 10.    ANSWER